"If Ben invests $4000 at 4% interest per year, how much additional money must he invest at 5 1/2% annual interest to ensure that the interest he receives each year is 4 1/2 % of the total amount invested?"

I really just don't understand this question at all! I tried the (P)(R)(T) = I formula and am just stumped!
Any help would be really appreciated!
L.Nytz

Answer by stanbon(57377)

"If Ben invests $4000 at 4% interest per year, how much additional money must he invest at 5 1/2% annual interest to ensure that the interest he receives each year is 4 1/2 % of the total amount invested?"
----------------
Let "x" be the amount of additional money he needs to invest.
EQUATION:
interest + interest = interest
0.04(4000) + 0.055x = 0.045(4000+x)
160 +0.055x = 180+0.045x
0.010x=20
x= $2000 (amount of additional money he needs to inverst at 5 1/2 %)
Cheers,
Stan H.